A bank teller has 54 $4 and $20 bills in her cash drawer. The value of the bills of $780. How many $5 bills are there?
I tried X+Y= 54
X-Y= 20
2X=34
X-17 $5 BILLS
THIS DOES NOT SEEM RIGHT

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

A bank teller has 54 $5 and $20 bills in her cash drawer. The value of the bills of $780. How many $5 bills are there?
Quantity Eq: f + t = 54 bills
Value Eq: 5f+20t = 780 dollars
-----
Divide thru the Value Eq by 5
f+t = 54
f+4t=156
---
Subtract and solve for "t":
3t =102
t = 34 (# of $20 bills)
---
Solve for "f":
f+t = 54
f+34 = 54
f = 20 (# of $5 bills)
======
Cheers,
Stan H.
===========

Answer by mananth(16946)

$5 bills x
$20 bill ------------ y
x+y= 54 .............1
5x+20y =780 .............2

multiply (1)by -20
Multiply (2) by 1
-20x-20y =-1080
5x+20y=780
Add the two equations
-15x=-300
/-15
x=20 $5 bills
plug value of x in (1)

x+y=54
20+y=54
y=54-20
y=34
y=34 $20 bill